LOFAR Newsletters December 2016

Distributed on 22 December 2016




-       The Conference ‘The Broad Impact of Low Frequency observing' will take place between 19-23 June 2017 in Bologna (Italy). It will run from Monday morning, 9 am till Friday, lunchtime. On Friday afternoon, we will held the 4th LOFAR Users Meeting at the same venue. The registration will open with the second announcement, which will be distributed shortly.

-       At the Cycle 7 PC meeting, 34 proposals (including the already active long-term ones) were awarded in total 1750 observing and 1478 processing hours, respectively.

-       Since end of November 2016 CEP4 is used as the main production cluster. Some issues remain with data losses for specific observations and bugs in the Pulsar pipelines. These are being worked on.

-       On 1 January 2017 CEP2 will be switched off and removed from his current location.

-       VIRAC (Latvia) plans to build a LOFAR station at its Irbene observatory. The Latvian ministry and parliament have already allocated the first tranche of money, for 2017. VIRAC is now working with Astrotec and ASTRON to prepare a 3-year implementation plan for 2017-2019. VIRAC plans to be an early adopter of LOFAR2.0 technology, for which a system design effort is getting under way right now at ASTRON.



Array status:


-       38 stations operational in the Netherlands: 24 core stations and 14 remote stations.

-       12 international stations operational: DE601, DE602, DE603, DE604, DE605, FR606, SE607, UK608, DE609, PL610, PL611, PL612

-       A new LOFAR station (IE613) is being built on the grounds of Birr Castle, located centrally in Ireland. Expected completion: first half 2017.

-       The LCU's of all Dutch stations were replaced on 15-16 November 2016. The replacement for the international stations completed on 15 December 2016.

-       The overview of non-operational antenna elements for LBA and HBA is available here.

-       Station calibration:

  • Data collection in one SPARSE mode will be performed during the Christmas holidays for the whole Dutch array. Previous runs performed only on the remote stations have shown that the updated calibration software is now able to manage SPARSE data.
  • New calibration tables in HBA LOW and HIGH will be installed by the end of the month.



Observing System Status:


-       Nominal system performance during the period November - December 2016. Debugging activities associated with CEP4 commissioning lowered the observing efficiency during this period.



CITT Status (E. Orru' & T. J. Dijkema)


  • Pre-factor pipeline: work is on going in order to allow ingesting of the output products into the LTA and to speed up the Initial Subtract step.
  • FACTOR pipeline: the code was heavily updated to use WSclean for imaging and NDPPP for solving. Commissioning and bug fixing are on going.
  • WSclean run with IDG. The run was successful but a bug was identified - work to solve this is on going.
  • LBA: a preliminary version of a calibrator pipeline is available for
    testing on github (Pill).
  • A direction dependent error (DDE) solver is being implemented in DPPP with the ultimate goal to perform calibration at higher time and frequency resolutions.



MSSS Status (J. Broderick & G. Heald)



-       Progress continues towards finalizing the reprocessing strategy for a higher-resolution (approx. 45 arcsec) HBA catalog. Our testing indicates that the full-bandwidth mosaics have noise levels as good as ~3 mJy/beam, and ~10 mJy/beam for each six-field, single-band mosaic. A recent review of the MSSS projects and working groups has also taken place, not only for planning purposes related to the aforementioned catalog, but also to steer efforts towards maximizing the scientific yield from the current MSSS-HBA internal data release (catalog v0.3). Moreover, an updated MSSS publication policy is available on the LOFAR wiki.



Observing Programmes



-       Cycle 7 observing programme: 15% complete. The observing schedule can be found here

-       Cycle 6 observing programme: 86% complete. The rest is being observed with second priority during Cycle 7.



CEP news:


-       CEP4

  • Characterization of the various RO processing pipelines is on going.

-      CEP2:

  • The cluster will be switched off as per 1 January 2017.

-       CEP3:

  • Cluster info and schedule available here.
  • Monitoring of node usage is active. Statistics will be used to evaluate requests for extensions.



Calendar next LOFAR activities:


Note: the following events are now marked on an online calendar that is available here.


-       ‘The Broad Impact of Low Frequency observing': 19-23 June 2017, Bologna (Italy). 

-       ‘Long baseline Busy Week': 9-13 January 2017, ASTRON

-       Next LSM's : 11/01, 08/02

(all presentations given at the LSM and video recordings are available here).

-       Next Stop days: 03/01, 07/02

-       Next software roll outs: 30/01, 13/03

-       Next LOFAR bulletin: February 2017

Design: Kuenst.    Development: Dripl.    © 2022 ASTRON